OpenStudy (kitkat1):

Medal for best answer!!!!! Given the system of equations presented here: 4x + y = 4 2x + 7y = 28 Which of the following actions creates an equivalent system such that, when combined with the other equation, one of the variables is eliminated? Multiply the second equation by −1 to get −2x − 7y = −28 Multiply the second equation by −4 to get −8x − 28y = −112 Multiply the first equation by −7 to get −28x − 7y = −28 Multiply the first equation by −2 to get −8x − 2y = −8

OpenStudy (kitkat1):

@whpalmer4 can u help me?

OpenStudy (kitkat1):

Does anybody know this?

OpenStudy (danjs):

you want to multiply one of the equations by some constant value, so that when you add them together, one variable will add to zero

OpenStudy (kitkat1):

Multiply by -2?

OpenStudy (danjs):

notice if you multiply the second by (-2), you get 4x + y = 4 -4x -14 y = -56 -----------------

OpenStudy (danjs):

then when you add the two equations, the x term will drop out (-4 + 4) = 0x

OpenStudy (danjs):

you see that

OpenStudy (kitkat1):

Yes I do so the right answer is D?

OpenStudy (danjs):

no, that example i just did multiplied the second by -2

OpenStudy (kitkat1):

Okay but how do I find the answer?

OpenStudy (danjs):

so, instead, how would you make the Y term drop out

OpenStudy (kitkat1):

Multiply by -1?

OpenStudy (danjs):

try -7 times the first equation (-7)(4x + y = 4 ) 2x + 7y = 28

OpenStudy (danjs):

gives -28x - 7y = -28 2x + 7y = 28 ------------------

OpenStudy (danjs):

adding equation 1 to equation 2 now, the y term goes to zero -26x + 0y = 0

OpenStudy (danjs):

so -26x = 0 x=0 plug into either to get the y

OpenStudy (danjs):

x=0 , y=4 to solve systems of linear eqn you can - muyltiply equations by constants add equations together switch order of equations

OpenStudy (kitkat1):

I ended up with 43=28 when plugged into the 2nd equation I don't know If I done it wrong.

OpenStudy (kitkat1):

I dont know what I am doing that I dont get the answer

OpenStudy (danjs):

start system 4x + y = 4 2x + 7y = 28 --------------- multiply first equation through by -7 to get -28x - 7 y = -28 2x + 7y = 28 -------------------

OpenStudy (kitkat1):

So I plug -7 to each of those equations the positive and the negative one?

OpenStudy (danjs):

adding the equations together eliminates y, what the instructions say to do -28x + 2x - 7y + 7y = -28 + 28 -26x =0

OpenStudy (danjs):

no, multiply each term in the first equation by a -7

OpenStudy (kitkat1):

Okay so the answer is 3?

OpenStudy (danjs):

multiply first equation through by -7 to get , second equation is not changed -28x - 7 y = -28 2x + 7y = 28 -------------------

OpenStudy (kitkat1):

Yes so the 3rd choice

OpenStudy (danjs):

th epoint of that was to eliminate the y variable when you add the two equations together

OpenStudy (danjs):

-26x = 0

OpenStudy (danjs):

so x=0

OpenStudy (danjs):

and if you want to find y, plug x=0 into either of the equations

OpenStudy (danjs):

4x + y = 4 2x + 7y = 28

OpenStudy (danjs):

y = 4, should be the same in both , or you messed up

OpenStudy (kitkat1):

Yes I got that so 4=4

OpenStudy (danjs):

x = 0, so 4(0) + y = 4 ----->y = 4 2(0) + 7y = 28 -----7y = 28 ---->y = 4 same y=4 from both equations when x=0

OpenStudy (kitkat1):

Okay I dont get whats the answer should be

OpenStudy (danjs):

another answer they dont include, was what i did at the very start, i figured y first, multiplied second equation by -2, so when added the x term goes away

OpenStudy (kitkat1):

So which option should I pick then?

OpenStudy (danjs):

The third one, we multiplied the fisrt by -7 to get that Which of the following actions creates an equivalent system such that, when combined with the other equation, one of the variables is eliminated? -when combined means when you add them together, one variable will go to zero, like the -7y + 7y = 0y , y is eliminated

OpenStudy (kitkat1):

Okay can you help me with one more question?

OpenStudy (danjs):

just takes a few practice probs to get it down,

OpenStudy (danjs):

sure i can do one more

OpenStudy (kitkat1):

Determine the solution to f(x) = g(x) using the following system of equations f(x) = 5.5x − 13 g(x) = −5x + 18.5 x = 1 x = 2 x = 3 x = 4

OpenStudy (kitkat1):

I remember doing this but cant remember how

OpenStudy (danjs):

both f(x) and g(x) are functions on the XY plane. when f(x) = g(x), the two functions intersect at the same point (x,y)

OpenStudy (kitkat1):

Can u show me how to solve it?

OpenStudy (danjs):

set them equal f(x) = g(x) 5.5x - 13 = -5x + 18.5

OpenStudy (kitkat1):

Then?

OpenStudy (danjs):

add 5x to both sides

OpenStudy (kitkat1):

like this 5.5x - 13 - -5x +18.5 5x 5x

OpenStudy (danjs):

10.5x - 13 = 18.5 add 13 both sides 10.5x = 31.5

OpenStudy (danjs):

divide both sides by 10.5 1x = 31.5 / 10.5

OpenStudy (kitkat1):

Okay then divide by 10.5

OpenStudy (danjs):

x = 3

OpenStudy (danjs):

yep, the goal is to isolate x by itself

OpenStudy (kitkat1):

Okay thank you very much for the help!

OpenStudy (danjs):

no prob, anytime